OpenStudy (anonymous):

A car travels 172 miles using 7 gallons of gas. At that rate, how far can the car travel using 42 gallons of gas?

OpenStudy (anonymous):

\[\frac{172}{7} = \frac{ x }{ 42 }\] Set up the above ratio and solve for "x". Staying away from decimals and simplifying, you get: \[\frac{ 172 \times 42 } { 7 } = x = 172 \times 6\]